OpenStudy (anonymous):

How can i solve the quadratic equation x2 - 4x + 57 = -5 by completing the square?

OpenStudy (johnweldon1993):

First things first...you want it in the form \[\large ax^2 + bx = c\] so subtract 57 from both sides of the equation...what do we have now?

OpenStudy (anonymous):

x2 - 4x = -5 ?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Remember you have to subtract it from the other side too... \[\large x^2 - 4x + 57 - 57 = -5 - 57\] what do you have after that?

OpenStudy (anonymous):

dont know /: @johnweldon1993

OpenStudy (johnweldon1993):

Don't get discouraged...*ps...this is going to be a complex answer I can already tell... \[\large x^2 - 4x + 57 - 57 = -5 - 57\] Just simplify here.... + 57 - 57 = 0....and -5 - 57 = -62...so now we have \[\large x^2 - 4x = -62\] okay?

OpenStudy (anonymous):

Just wrote that down, thank you! @johnweldon1993

OpenStudy (johnweldon1993):

We are not even CLOSE to done I hope you know that lol...

OpenStudy (johnweldon1993):

That was just getting it into the right form to do the procedure....now we begin :D ready?

OpenStudy (anonymous):

Oh i thought we were lol awks. Yes :) @johnweldon1993

OpenStudy (johnweldon1993):

Nope :P lol okay...now you have an equation in the form \[\large ax^2 + bx = c\] *compare it to your equation... \[\large x^2 - 4x = -62\] look the same?

OpenStudy (anonymous):

Yes thats the standard form of a quadratic functio, that's what i was trying to find lol @johnweldon1993 totally forgot how it was formed

OpenStudy (anonymous):

*function

OpenStudy (johnweldon1993):

Perfect....now have you done 'completing the square' before?

OpenStudy (anonymous):

Can you refresh my mind on what it is again? @johnweldon1993

OpenStudy (johnweldon1993):

Absolutely :) So you get you equation into standard form *check* lol Now you take your 'b' value.....divide it by 2....then square the result...what is this?

OpenStudy (anonymous):

2? @johnweldon1993

OpenStudy (johnweldon1993):

Not quite...you didnt square it... your 'b' value is -4 right ....you divide that by 2.....-4/2 = -2.......now you square that result....-2² = 4....right?

OpenStudy (anonymous):

Oh yes, my mistake @johnweldon1993

OpenStudy (johnweldon1993):

No problem...so now...you add that to both sides of the equation... \[\large x^2 - 4x + 4 = -62 + 4\] simplified down a bit we have \[\large x^2 - 4x + 4 = -58 \] Now you would factor the left hand side...but...remember what I said...this is going to have imaginary results...*complex numbers* can you do that?

OpenStudy (johnweldon1993):

We know we are looking for a perfect square here....so what number times itself...= 4....but also when added to itself = -4?

OpenStudy (anonymous):

-2 @johnweldon1993

OpenStudy (johnweldon1993):

Right...so we have \[\large (x - 2)^2 = -58\] Now take the square root of both sides *this is where the complex numbers comes in* lol...and I'll brb have an errand to run :)

OpenStudy (johnweldon1993):

I hope you got this @justagal1619 From the step before....take the aquare root of (x - 2)^2 = -58 this becomes \[\large \sqrt{(x - 2)^2} = \sqrt{-58}\] Which simplifies into \[\large (x - 2) = \pm \sqrt{-58}\] Now when you solve for 'x'...you just add 2 to both sides of the equation... \[x = 2 \pm \sqrt{-58} \]

OpenStudy (anonymous):

@johnweldon1993 kinda thanks anyways :-)